# Two-Channel Filter Banks on Joint Time-Vertex Graphs with Oversampled Graph Laplacian Matrix

### v1

@@ I Introduction
WITH the rapid advancement of information technology, the global volume of stored and processed data has been growing exponentially. A significant portion of these data are generated on complex networks with inherently irregular structures, which are naturally modeled as graphs, while the associated data are represented as time-varying graph signals. In recent years, graph signal processing (GSP) has evolved rapidly, establishing itself as a core framework for analyzing and processing such data [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
@@
Graphs serve as a powerful tool for representing irregular and complex data and have found widespread applications in social networks, sensor networks, transportation systems, as well as image and video processing [13, 14, 15]. Within this paradigm, signals are mapped to graph vertices, enabling GSP to provide effective tools for network data processing. Considerable efforts have been devoted to extending classical time–frequency methods to the graph domain, including discrete Fourier transform (DFT) [1, 2, 3], windowed Fourier transform [6], spectral analysis [7], filtering [8], sampling and interpolation [9, 10], and uncertainty principles [10, 11]. Analogous to the role of DFT [12] in conventional signal processing, the graph Fourier transform (GFT) has become fundamental to many GSP applications, such as denoising [16, 17, 18, 19], reconstruction [21], classification [9], clustering [22, 23, 24, 25], segmentation [26], estimation [27, 28], semi-supervised learning [29], graph learning [30, 31], and the development of graph neural networks [32, 33, 34].
@@
However, many graph-structured data also evolve over time. For instance, video sequences can be viewed as temporally evolving images, and user behaviors in social networks often vary dynamically. This motivates the need for a unified framework for analyzing joint time–vertex signals. Pioneering works [22, 35] integrated discrete-time processing with GSP, laying the foundation for this direction. The joint time–vertex Fourier transform (JFT), which combines the DFT in the temporal domain and the GFT in the vertex domain, serves as a key tool for joint spectral analysis. Building on this framework, a variety of applications have emerged, including reconstruction of time-varying graph signals and images [36, 37, 38, 39], prediction of joint spectral–temporal information [40], and modeling of stationary dynamics on graphs [41]. Spatio-temporal image and video restoration can be regarded as a special case of joint time–vertex signals, further highlighting the importance of developing joint transforms and filtering methods [22, 42].
@@
To further enhance the expressive power of joint time–vertex signal processing, multiresolution analysis plays an important role. Multiresolution analysis has long been effective in traditional analysis, compression, and denoising [43], and its extension to the joint domain provides a powerful tool for decomposition and representation. In GSP, two-channel critically sampled wavelet filter banks have been proposed to realize multiresolution transforms [16, 17]. For bipartite graphs, such filter banks avoid spectral folding during downsampling and ensure perfect reconstruction [18]. However, for arbitrary non-bipartite graphs, the original graph must be partitioned into bipartite subgraphs, each retaining all vertices but only a subset of edges, whose union reconstructs the original graph. Applying filter banks to each subgraph leads to multi-dimensional transforms, yet many original edges are not fully utilized in a single stage. To address this limitation, oversampled graph Laplacian matrices were introduced in [18, 19, 20], enabling flexible oversampling through the addition of nodes and edges, and thereby supporting redundant signal representations. Notably, the filter bank concept has proven effective for graph-based decomposition and reconstruction and has inspired graph-based image denoising.
@@
Image denoising aims to remove noise while preserving structural details, and GSP has attracted significant attention for this purpose [44]. Early studies employed graph Laplacian regularization to enhance smoothness while retaining edges, and subsequent extensions with manifold-based gradient regularizers and plug-and-play unrolled optimization approaches further improved denoising performance and flexibility [45, 46, 47]. Multiresolution and reversible graph Fourier transforms were developed for denoising, as well as image compression and watermarking [48, 49]. More recent work has applied graph convolutional techniques to challenging denoising tasks, including point cloud denoising, dequantization, and video super-resolution [50, 51, 52, 53]. In parallel, graph-based wavelet and filter bank methods have been proposed for image modeling. For example, graph wavelet filter banks provide sparse, edge-aware image representations, achieving superior performance over conventional wavelets in nonlinear approximation and denoising [16, 17]. The oversampling concept further enhances these filter banks, improving robustness in image denoising and reconstruction [18, 19, 20]. These works demonstrate the strong potential of GSP in modern image and video denoising and restoration, motivating extensions to joint time–vertex graph methodologies.
@@
Against this background, this paper develops a unified framework for joint time–vertex signal processing. The framework is built on an oversampled graph Laplacian matrix extended from the vertex domain to the temporal domain, supporting bipartite expansion and redundant multiresolution representations. Based on this operator, a two-channel joint time–vertex oversampled filter bank is designed to satisfy perfect reconstruction on bipartite graphs and realize oversampling in both time and vertex domains via a $K$-coloring strategy. The framework is applied to signal and image/video denoising, with experiments demonstrating improved decomposition and reconstruction performance compared to existing methods, highlighting its potential at the intersection of GSP and image processing.
@@
  • Introduce a joint time–vertex oversampled graph Laplacian matrixthat extends vertex domain oversampling to the temporal domain, preserving the original joint graph structure while enabling bipartite expansion and redundant multiresolution representations.
  • Design a two-channel joint time–vertex oversampled filter bank that satisfies perfect reconstruction conditions on bipartite graphs and implements oversampling in both time and vertex domains using a $K$-coloring strategy.
  • Apply the proposed framework to joint time–vertex signal and image/video denoising, demonstrating its effectiveness in decomposition and reconstruction, and outperforming existing methods in numerical experiments.
The remainder of the paper is organized as follows. Section II reviews joint time–vertex signal processing, two-channel oversampled graph filter banks, and oversampled graph Laplacians. Section III introduces the proposed time oversampled Laplacian and joint oversampled Laplacian, together with their properties. Section IV presents the design of critically sampled and oversampled joint filter banks. Section V discusses bipartite extensions in both time and vertex domains. Section VI reports experimental results on signal decomposition, reconstruction, and denoising, including image and video applications. Finally, Section VII concludes the paper.
@@ II Preliminaries
A brief overview of existing theories on joint time–vertex signal processing [22], graph filter banks, and oversampled graph Laplacian matrix [18, 19] is provided below.
@@ II-A Joint Time-Vertex Signal Processing
Both $\mathbf{X}$ and its vectorized form $\bm{x}=\text{vec}(\mathbf{X})\in\mathbb{R}^{NT}$ are referred to as joint time–vertex signals [22]. In $\mathbf{X}$, each column corresponds to the state of the graph at a specific time, while each row describes the temporal evolution of a single vertex. The GFT [1, 2, 3] along the vertex dimension is defined as
@@ III-A Oversampled Time Laplacian Matrix
This operator retains the desirable spectral properties of the original time Laplacian and simultaneously plays a central role in constructing the joint time–vertex oversampled Laplacian.
@@ III-B Oversampled Time-Vertex Laplacian Matrix
Building upon the oversampled graph constructions in both the time and vertex domains, we now propose a joint time–vertex oversampled graph Laplacian to support unified modeling and filter bank design across both dimensions.
@@
This formulation reveals that the joint spectrum is composed by summing all possible pairs of temporal and vertex eigenvalues, while the joint eigenvectors are constructed as the Kronecker product of their respective bases. As a result, joint filtering operations can be efficiently implemented as separable filters acting independently across the time and vertex dimensions.
To align with the joint time-vertex oversampled graph structure $\widetilde{\mathcal{T}}\otimes\widetilde{\mathcal{G}}$, the original time–vertex signal must be accordingly extended. Let the original signal be denoted as $\mathbf{X}_{0,0}\in\mathbb{R}^{N_{0}\times T_{0}}$, defined on the original product graph $\mathcal{T}\otimes\mathcal{G}$, where $T_{0}$ and $N_{0}$ are the numbers of temporal and vertex nodes, respectively.
@@ IV-A Joint Critically Sampled Graph Filter Banks
The time–vertex graph $\mathcal{J}=\mathcal{T}\otimes\mathcal{G}$ enables joint processing along both domains. The input signal $\mathbf{X}$ is first filtered in the vertex and temporal domains using analysis filters $\mathbf{H}^{k}_{\mathcal{G}}$ and $\mathbf{H}^{k}_{\mathcal{T}}$, where $\mathbf{H}^{k}_{\mathcal{G}}$ is defined as Eq. (5), and another is
@@ IV-B Joint Oversampled Graph Filter Banks
To enable flexible signal modeling and transformation in the joint time–vertex domain, we introduce two-channel oversampled graph filter banks. Unlike the critically sampled case [17], the oversampled structure preserves all nodes in both time and vertex domains, allowing richer representation and enabling integration into learning frameworks [19].
@@ V-A Construction of Time Bipartite Graph
Within the joint time–vertex signal processing framework, we propose a method for constructing a bipartite graph structure in the time domain. The key idea is to model the temporal sequence as a ring graph and employ graph coloring theory to design an oversampled bipartite structure. The detailed construction is presented in Appendix A.
@@
V-B Joint Extension of the Time–Vertex Domain
To support the design of joint filter banks, we propose a novel method for constructing oversampled bipartite graphs in the joint time–vertex domain based on graph coloring strategies. This method independently builds oversampled bipartite structures in both the time and vertex domains, and subsequently combines them via a tensor product. The resulting graph preserves all original adjacency relations while enabling flexible control of redundancy, thereby facilitating efficient joint signal processing.
@@ VI Experiments and Results
This section presents experiments on arbitrary graphs, images, and videos, employing the Harary algorithm [54, 55] to evaluate the performance gains of the joint time–vertex oversampled Laplacian in reconstruction and denoising.
@@ VI-A Reconstruction of Joint Time-Vertex Graph Signal
To evaluate the reconstruction performance of the proposed oversampled joint time–vertex filter bank, we conduct epidemic spreading simulations on the Minnesota Road Network [56]. The graph consists of $N=2642$ nodes, each representing an intersection, with edges denoting road connections, and is widely adopted for modeling diffusion processes related to mobility and transportation. The epidemic dynamics are modeled using the Susceptible–Exposed–Infectious–Recovered–Susceptible (SEIRS) framework [22, 23, 24], where recovered individuals lose immunity after a finite period. The state variables are denoted as $S(t)$, $E(t)$, $I(t)$, and $R(t)$, representing the numbers of susceptible, exposed, infectious, and temporarily immune individuals, respectively. The discrete-time evolution is governed by
@@ VI-B Joint Time-Vertex Graph Signal Denoising
To provide further insights, Fig. 6 and Fig. 7 present qualitative comparisons of denoising results for the Yale Coat of Arms and the Minnesota Road Network under Gaussian noise with $\sigma=1/2$, respectively. Overall, the oversampled filter bank exhibits clear advantages in joint time-vertex signal denoising, highlighting its potential in complex signal processing scenarios.
@@ VI-C Image Denoising via Joint Oversampled Filter Banks
To evaluate the efficacy of the proposed two-channel joint time–vertex oversampled graph filter banks, we conduct experiments on both static images and dynamic videos. Within the joint time–vertex framework, an image or a sequence of frames is modeled as a graph, where pixels are treated as vertices and temporal evolution is captured along the time dimension. This allows simultaneous filtering across spatial and temporal domains.
@@
Results are shown in Figs. 8 and 9, where panels (a) and (e) correspond to the original images, (b) and (f) to the noisy inputs, (c) and (g) to reconstructions using critically sampled filter banks, and (d) and (h) to reconstructions using oversampled filter banks. Visually, the oversampled method more effectively suppresses noise while preserving edges and fine details. Quantitative evaluations under different noise levels are summarized in Tables IV and V, which report SNR improvements for both image pairs.
@@
For dynamic video denoising, the walking sequence from the human actions video database222[Online]. Available: https://www.csc.kth.se/cvap/actions/ is used. The first 3 seconds (approximately 75 frames) are extracted, converted to grayscale, and resized to $N\times N$, forming a spatial graph with each pixel as a node. The temporal dimension is represented by a ring graph over the frame sequence, yielding a joint time-vertex graph structure. Gaussian noise with standard deviation $\sigma=10$ is added to the frames, and both oversampled and critically sampled filter banks are applied for denoising. Selected key frames are plotted in Fig. 10, with rows corresponding to original frames, noisy frames, critically sampled reconstructions, and oversampled reconstructions. Quantitatively, the noisy video exhibits an SNR of 22.43 dB, which improves to 23.77 dB with the critically sampled filter bank and further to 25.85 dB with the oversampled filter bank. The oversampled filter bank effectively suppresses noise along both spatial and temporal dimensions while preserving motion details and temporal continuity, producing smoother and more natural motion trajectories.

### v2

@@ I Introduction
WITH the rapid advancement of information technology, the global volume of stored and processed data has been growing exponentially. A significant portion of these data are generated on complex networks with inherently irregular structures, which are naturally modeled as graphs, while the associated data are represented as time-varying graph signals. In recent years, graph signal processing (GSP) has evolved rapidly, establishing itself as a core framework for analyzing and processing such data [1, 2, 3, 5, 4, 6, 7, 8, 9, 10, 11, 12].
@@
Graphs serve as a powerful tool for representing irregular and complex data and have found widespread applications in social networks, sensor networks, transportation systems, as well as image and video processing [14, 15, 16]. Within this paradigm, signals are mapped to graph vertices, enabling GSP to provide effective tools for network data processing. Considerable efforts have been devoted to extending classical time-frequency methods to the graph domain, including discrete Fourier transform (DFT) [1, 2, 3, 4], windowed Fourier transform [7], spectral analysis [8], filtering [9], sampling and interpolation [10, 11], and uncertainty principles [11, 12]. Analogous to the role of DFT [13] in conventional signal processing, the graph Fourier transform (GFT) has become fundamental to many GSP applications, such as denoising [17, 18, 19, 20], reconstruction [22], classification [10], clustering [23, 24, 25, 26], segmentation [27], estimation [28, 29], detection [30], semi-supervised learning [31], graph learning [32, 33], and the development of graph neural networks [34, 35, 36].
@@
However, many graph-structured data also evolve over time. For instance, video sequences can be viewed as temporally evolving images, and user behaviors in social networks often vary dynamically. This motivates the need for a unified framework for analyzing joint time-vertex signals. Pioneering works [23, 37] integrated discrete-time processing with GSP, laying the foundation for this direction. The joint time-vertex Fourier transform (JFT), which combines the DFT in the temporal domain and the GFT in the vertex domain, serves as a key tool for joint spectral analysis. Building on this framework, a variety of applications have emerged, including reconstruction of time-varying graph signals [38, 39, 40, 41], prediction of joint spectral-temporal information [42], and modeling of stationary dynamics on graphs [43]. Spatio-temporal image and video restoration can be regarded as a special case of joint time-vertex signals, further highlighting the importance of developing joint transforms and filtering methods [23, 44].
@@
To further enhance the expressive power of joint time-vertex signal processing, multiresolution analysis plays an important role. Multiresolution analysis has long been effective in traditional analysis, compression, and denoising [45], and its extension to the joint domain provides a powerful tool for decomposition and representation. In GSP, two-channel critically sampled wavelet filter banks have been proposed to realize multiresolution transforms [17, 18]. For bipartite graphs, such filter banks avoid spectral folding during downsampling and ensure perfect reconstruction [19]. However, for arbitrary non-bipartite graphs, the original graph must be partitioned into bipartite subgraphs, each retaining all vertices but only a subset of edges, whose union reconstructs the original graph. Applying filter banks to each subgraph leads to multi-dimensional transforms, yet many original edges are not fully utilized in a single stage. To address this limitation, oversampled graph Laplacian matrices were introduced in [19, 20, 21], enabling flexible oversampling through the addition of nodes and edges, and thereby supporting redundant signal representations. Notably, the filter bank concept has proven effective for graph-based decomposition and reconstruction and has inspired graph-based image denoising.
@@
Image denoising aims to remove noise while preserving structural details, and GSP has attracted significant attention for this purpose [46]. Early studies employed graph Laplacian regularization to enhance smoothness while retaining edges, and subsequent extensions with manifold-based gradient regularizers and plug-and-play unrolled optimization approaches further improved denoising performance and flexibility [47, 48, 49]. Multiresolution and reversible graph Fourier transforms were developed for denoising, as well as image compression and watermarking [50, 51]. In parallel, graph-based wavelet and filter bank methods have been proposed for image modeling. For example, graph wavelet filter banks provide sparse, edge-aware image representations, achieving superior performance over conventional wavelets in nonlinear approximation and denoising [17, 18]. The oversampling concept further enhances these filter banks, improving robustness in image denoising and reconstruction [19, 20, 21]. These works demonstrate the strong potential of GSP in modern image and video denoising and restoration, motivating extensions to joint time-vertex graph methodologies.
@@
Against this background, this paper develops a unified framework for joint time-vertex signal processing. The framework is built on an oversampled graph Laplacian matrix extended from the vertex domain to the temporal domain, supporting bipartite expansion and redundant multiresolution representations. Based on this operator, a two-channel joint time-vertex oversampled filter bank is designed to satisfy perfect reconstruction on bipartite graphs and realize oversampling in both time and vertex domains via a $K$-coloring strategy. The framework is applied to signal and image/video denoising, with experiments demonstrating improved decomposition and reconstruction performance compared to existing methods, highlighting its potential at the intersection of GSP and image processing.
@@
  • Introduce a joint time-vertex oversampled graph Laplacian matrixthat extends vertex domain oversampling to the temporal domain, preserving the original joint graph structure while enabling bipartite expansion and redundant multiresolution representations.
  • Design a two-channel joint time-vertex oversampled filter bank that satisfies perfect reconstruction conditions on bipartite graphs and implements oversampling in both time and vertex domains using a $K$-coloring strategy.
  • Apply the proposed framework to joint time-vertex signal and image/video denoising, demonstrating its effectiveness in decomposition and reconstruction, and outperforming existing methods in numerical experiments.
The remainder of the paper is organized as follows. Section II reviews joint time-vertex signal processing, two-channel oversampled graph filter banks, and oversampled graph Laplacians. Section III introduces the proposed time oversampled Laplacian and joint oversampled Laplacian, together with their properties. Section IV presents the design of critically sampled and oversampled joint filter banks. Section V discusses bipartite extensions in both time and vertex domains. Section VI reports experimental results on signal decomposition, reconstruction, and denoising, including image and video applications. Finally, Section VII concludes the paper.
@@ II Preliminaries
A brief overview of existing theories on joint time-vertex signal processing [23], graph filter banks, and oversampled graph Laplacian matrix [19, 20] is provided below.
@@ II-A Joint Time-Vertex Signal Processing
Both $\mathbf{X}$ and its vectorized form $\bm{x}=\text{vec}(\mathbf{X})\in\mathbb{R}^{NT}$ are referred to as joint time-vertex signals [23]. In $\mathbf{X}$, each column corresponds to the state of the graph at a specific time, while each row describes the temporal evolution of a single vertex. The GFT [1, 2, 3] along the vertex dimension is defined as
@@ III-A Oversampled Time Laplacian Matrix
This operator retains the desirable spectral properties of the original time Laplacian and simultaneously plays a central role in constructing the joint time-vertex oversampled Laplacian.
@@ III-B Oversampled Time-Vertex Laplacian Matrix
Building upon the oversampled graph constructions in both the time and vertex domains, we now propose a joint time-vertex oversampled graph Laplacian to support unified modeling and filter bank design across both dimensions.
@@
This construction guarantees that the spectral structure of the joint Laplacian remains separable. Although Eq. (11) is structurally similar to the separable joint Laplacian operators [23] defined by Eq. (3), our framework does not operate directly on the original product graph. Instead, both the time and vertex domains are first extended through structured oversampling. Therefore, the resulting joint Laplacian is defined on an augmented graph that supports a redundant two-channel filter bank design, rather than a purely separable joint filtering framework.
@@
This formulation reveals that the joint spectrum is composed by summing all possible pairs of temporal and vertex eigenvalues, while the joint eigenvectors are constructed as the Kronecker product of their respective bases. This separable spectral structure facilitates efficient implementation of the proposed joint oversampled filter bank, while the redundancy introduced by oversampling fundamentally distinguishes our framework from conventional separable joint filtering methods.
To align with the joint time-vertex oversampled graph structure $\widetilde{\mathcal{T}}\otimes\widetilde{\mathcal{G}}$, the original time-vertex signal must be accordingly extended. Let the original signal be denoted as $\mathbf{X}_{0,0}\in\mathbb{R}^{N_{0}\times T_{0}}$, defined on the original product graph $\mathcal{T}\otimes\mathcal{G}$, where $T_{0}$ and $N_{0}$ are the numbers of temporal and vertex nodes, respectively.
@@
Remark 2: Unlike conventional joint filtering on the original product graph, the proposed transform underpins a redundant joint time-vertex filter bank, enabling multiresolution representation and enhanced reconstruction stability.
@@ IV-A Joint Critically Sampled Graph Filter Banks
The time-vertex graph $\mathcal{J}=\mathcal{T}\otimes\mathcal{G}$ enables joint processing along both domains. The input signal $\mathbf{X}$ is first filtered in the vertex and temporal domains using analysis filters $\mathbf{H}^{k}_{\mathcal{G}}$ and $\mathbf{H}^{k}_{\mathcal{T}}$, where $\mathbf{H}^{k}_{\mathcal{G}}$ is defined as Eq. (5), and another is
@@ IV-B Joint Oversampled Graph Filter Banks
To enable flexible signal modeling and transformation in the joint time-vertex domain, we introduce two-channel oversampled graph filter banks. Unlike the critically sampled case [18], the oversampled structure preserves all nodes in both time and vertex domains, allowing richer representation and enabling integration into learning frameworks [20].
The normalized Laplacian of a bipartite graph has a spectrum symmetric about 1, so if $\lambda$ is an eigenvalue, $2-\lambda$ is also an eigenvalue (Eq. (6)). This enables quadrature mirror kernels that cancel aliasing [19]. In the joint time-vertex setting, separability ensures that perfect reconstruction in each domain guarantees perfect reconstruction of the overall transform, extending classical graph filter bank to the joint domain.
@@ V-A Construction of Time Bipartite Graph
Within the joint time-vertex signal processing framework, we propose a method for constructing a bipartite graph structure in the time domain. The key idea is to model the temporal sequence as a ring graph and employ graph coloring theory to design an oversampled bipartite structure. The detailed construction is presented in Appendix A.
@@
V-B Joint Extension of the Time-Vertex Domain
To support the design of joint filter banks, we propose a novel method for constructing oversampled bipartite graphs in the joint time-vertex domain based on graph coloring strategies. This method independently builds oversampled bipartite structures in both the time and vertex domains, and subsequently combines them via a tensor product. The resulting graph preserves all original adjacency relations while enabling flexible control of redundancy, thereby facilitating efficient joint signal processing.
@@ V-B Joint Extension of the Time-Vertex Domain
By construction, $\rho_{\mathcal{T}}\geq 1$ and $\rho_{\mathcal{G}}\geq 1$, implying the lower bound $\rho\geq 1$, with equality attained when both domains are inherently bipartite and no duplication is required. In the time domain, when the ring length is odd, the proposed extension introduces at most a constant number of additional nodes, yielding $\rho_{\mathcal{T}}\to 1$ as $T\to\infty$. Consequently, for large-scale graphs, the overall redundancy remains bounded and asymptotically approaches that of the vertex extension.
It is important to emphasize that the proposed scheme does not seek combinatorial minimality of redundancy. Rather, under the constraint of single-stage bipartite decomposition and full preservation of original adjacency relations, it achieves near-minimal redundancy while enabling structurally consistent joint filter bank design.
@@ VI Experiments and Results
This section presents experiments on arbitrary graphs, images, and videos, employing the Harary algorithm [52, 53] to evaluate the performance gains of the joint time-vertex oversampled Laplacian in reconstruction and denoising.
@@ VI-A Reconstruction of Joint Time-Vertex Graph Signal
To evaluate the reconstruction performance of the proposed oversampled joint time-vertex filter bank, we conduct epidemic spreading simulations on the Minnesota Road Network [54]. The graph consists of $N=2642$ nodes, each representing an intersection, with edges denoting road connections, and is widely adopted for modeling diffusion processes related to mobility and transportation. The epidemic dynamics are modeled using the Susceptible-Exposed-Infectious-Recovered-Susceptible (SEIRS) framework [23, 24, 25], where recovered individuals lose immunity after a finite period. The state variables are denoted as $S(t)$, $E(t)$, $I(t)$, and $R(t)$, representing the numbers of susceptible, exposed, infectious, and temporarily immune individuals, respectively. The discrete-time evolution is governed by
@@ VI-B Joint Time-Vertex Graph Signal Denoising
To provide further insights, Fig. 6 and Fig. 7 present qualitative comparisons of the denoising results for the Yale Coat of Arms and the Minnesota Road Network under Gaussian noise with $\sigma=1/2$, respectively.
@@
Fig. 8 further reports the quantitative comparison of the denoising performance in terms of MSE for (a) the Yale Coat of Arms and (b) the Minnesota Road Network under Gaussian noise with $\sigma=1$. The results are evaluated over a range of threshold values defined as $\mathrm{threshold}=\kappa\sigma$, where $\kappa$ denotes the scaling coefficient. It can be observed that, when the threshold parameter is properly tuned, the proposed oversampled scheme consistently achieves the lowest reconstruction error among the compared methods. Overall, the oversampled filter bank exhibits clear advantages in joint time-vertex signal denoising, highlighting its strong potential for complex and large-scale signal processing applications.
Figure 8: MSE as a Function of the Threshold Scaling Coefficient $\kappa$: (a) Yale Coat of Arms; (b) Minnesota Road Network.
@@ VI-C Image Denoising via Joint Oversampled Filter Banks
To evaluate the efficacy of the proposed two-channel joint time-vertex oversampled graph filter banks, we conduct experiments on both static images and dynamic videos. Within the joint time-vertex framework, an image or a sequence of frames is modeled as a graph, where pixels are treated as vertices and temporal evolution is captured along the time dimension. This allows simultaneous filtering across spatial and temporal domains.
@@
Results are shown in Figs. 9 and 10, where panels (a) and (e) correspond to the original images, (b) and (f) to the noisy inputs, (c) and (g) to reconstructions using critically sampled filter banks, and (d) and (h) to reconstructions using oversampled filter banks. Visually, the oversampled method more effectively suppresses noise while preserving edges and fine details. Quantitative evaluations under different noise levels are summarized in Tables IV and V, which report SNR improvements for both image pairs.
@@
For dynamic video denoising, the walking sequence from the human actions video database222[Online]. Available: https://www.csc.kth.se/cvap/actions/ is used. The first 3 seconds (approximately 75 frames) are extracted, converted to grayscale, and resized to $N\times N$, forming a spatial graph with each pixel as a node. The temporal dimension is represented by a ring graph over the frame sequence, yielding a joint time-vertex graph structure. Gaussian noise with standard deviation $\sigma=10$ is added to the frames, and both oversampled and critically sampled filter banks are applied for denoising. Selected key frames are plotted in Fig. 11, with rows corresponding to original frames, noisy frames, critically sampled reconstructions, and oversampled reconstructions. Fig. 12 illustrates the influence of the threshold scaling coefficient $\kappa$ on the video denoising performance. It can be observed that the proposed oversampled scheme consistently achieves the highest SNR across the entire range of $\kappa$ values.
Figure 12: SNR as a function of the threshold scaling coefficient $\kappa$ for video denoising.
The noisy video yields an SNR of 22.43 dB, which increases to 24.55 dB with the critically sampled filter bank and further to 25.85 dB using the oversampled filter bank. The corresponding execution times are 0.79 s for the critically sampled scheme and 1.07 s for the oversampled scheme. The performance gain stems from the redundancy introduced by the oversampled filter bank, which enhances reconstruction accuracy and denoising capability while better preserving motion details and temporal continuity.
This improvement comes at the expense of higher computational cost compared to the critically sampled design. The runtime increases proportionally with the oversampling factor $\rho$, as well as with the graph size and temporal dimension. Nevertheless, for moderate problem scales, the computational overhead remains acceptable and is justified by the significant performance gain.